# Understanding High-Dimensional Bayesian Optimization

### v1

@@ A.1 Gaussian Processes
With a gradient-based approach, this is done by maximizing Eq. (9), which is usually multi-modal and difficult to optimize. The gradient of the marginal log-likelihood w.r.t. is given by
@@ A.2 Bayesian Optimization
Appendix B Additional Experiments
B.1 Ranking of Optimization Algorithms
@@
B.2 MLE Gradients for Real-World Experiments
@@
B.3 High-dimensional Benchmark Functions
@@
Figs. 14 and 15 show the OTSD and performance plots for the 100-dimensional Levy and Griewank functions. Fig. 16 shows the two-dimensional versions of the Levy, Griewank, and Schwefel benchmark functions.
@@
B.4 Hard Optimization Problems
@@
Appendix C Popular Benchmarks Seem Simpler Than Expected
@@
Figs. 19 and 20 further show that BO consistently evaluates a large share of the parameters at the border. Fig. 19 shows this for DSP by (Hvarfner et al., 2024) whereas Fig. 20 uses MLE as proposed in Section 4. The general trend is that, during the course of the optimization, increasingly many parameters are evaluated at the border, which is consistent with Fig. 18.

### v2

@@ A.1 Gaussian Processes
With a gradient-based approach, this is done by maximizing Eq. (9), which is usually multi-modal and difficult to optimize. The gradient of the marginal log-likelihood w.r.t. $\theta_{i}$ is given by
@@ A.2 Bayesian Optimization
Appendix B Additional Implementation Details
B.1 Implementation of RAASP
We perturb the top 5% observations using a normal distribution with $\sigma=10^{-3}$, truncated within $\mathcal{X}$. For $d\geq 20$, we also generate samples with only a subset of dimensions perturbed, each with a $\frac{20}{d}$ probability. Of $4m$ random samples, $2m$ are global samples from a scrambled Sobol sequence, $m$ are local samples around the top 5%, perturbing all dimensions, and $m$ are local samples around the top 5%, perturbing 20 dimensions on average if $d\geq 20$, or all dimensions if $d<20$. We call the $2m$ local samples the RAASP samples. The starting points for the gradient-based optimization of the AF are drawn from the $4m$ overall samples using Boltzmann sampling (Ament et al., 2024).
B.2 Optimization of the Acquisition Function
We use the LogEI AF (Ament et al., 2024) for its numerical stability. It is maximized by evaluating it on 512 scrambled Sobol points, then selecting five starting points via Boltzmann sampling for gradient-based optimization using L-BFGS-B, with up to 2000 iterations. The budget of 2000 iterations is rarely exhausted, as the optimizer typically converges much earlier (see Fig. 12).
Figure 12: Number of gradient updates for the AF optimization for MSR, and with and without RAASP sampling. RAASP sampling reduces the number of gradient updates.
This aligns with previous work claiming that EI attains maxima close to good observations (Ament et al., 2024; Hvarfner et al., 2024) and our observation that the starting points for the gradient-based AF maximizer predominantly originate from the RAASP samples.
Appendix C Additional Experiments
C.1 Ranking of Optimization Algorithms
@@
C.2 MLE Gradients for Real-World Experiments
@@
C.3 High-dimensional Benchmark Functions
@@
Figs. 15 and 16 show the OTSD and performance plots for the 100-dimensional Levy and Griewank functions. Fig. 17 shows the two-dimensional versions of the Levy, Griewank, and Schwefel benchmark functions.
@@
C.4 Hard Optimization Problems
@@
Appendix D Popular Benchmarks Seem Simpler Than Expected
@@
Figs. 20 and 21 further show that BO consistently evaluates a large share of the parameters at the border. Fig. 20 shows this for DSP by (Hvarfner et al., 2024) whereas Fig. 21 uses MLE as proposed in Section 4. The general trend is that, during the course of the optimization, increasingly many parameters are evaluated at the border, which is consistent with Fig. 19.